OpenStudy (anonymous):

Today, Suzy is 17 years older than Billy, and next year she will be two times as old as Billy. How old is Suzy now? How old is Billy?

OpenStudy (anonymous):

Let's let Billy's age equal b, and Suzy's age equal b + 17. Next year, Billy will be one year older. So Billy's age will be b + 1. Suzy's age will be b + 17 + 1, or b + 18. So Suzy's age = 2(Billy's age). b + 18 = 2(b + 1) b + 18 = 2b + 2 b + 16 = 2b b = 16 So Billy is 16 now. That means Suzy is 33.

OpenStudy (anonymous):

awesome thanks